OpenStudy (anonymous):

The length of an arc of a circle is 12cm. The corresponding sector area is 108cm^2. find the radius of the circle?

OpenStudy (anonymous):

Let A the area, R the radius, a the angle in radian and L the length of the arc. Then A = 1/2 a R² = 108 and L = a R = 12 a R² = 216 and a R = 12 thus R = 216/12 = 18 cm the angle is a = 12/18 = 2/3 rad

OpenStudy (anonymous):

rad* radius

OpenStudy (anonymous):

I dont undertand.

OpenStudy (anonymous):

could you try and dumb down your answer for me?

OpenStudy (anonymous):

ok lets try this and see if it works

OpenStudy (anonymous):



OpenStudy (anonymous):

arc lenght you have is 12, and in general the arc length \(L\) given the angle \(\theta\) and the radius \(r\) is \(L=\theta r\) in your case you know \(12=\theta r\)

OpenStudy (anonymous):

the area \(A\) is 108 and in general the area given \(\theta\) and \(r\) is \(A=\frac{1}{2}\theta r^2\) so you also know that \(108=\frac{1}{2}\theta r ^2\) or \(216=\theta r^2\)

OpenStudy (anonymous):

so we have two equations, \[ 216=\theta r^2\] and \[12=\theta r\] ok i stopped

OpenStudy (anonymous):

i was using \(\theta\) in terms of radians, but it makes no difference for your problem because you are not looking for \(\theta\) just \(r\)

OpenStudy (anonymous):

divide and you get \[\frac{216}{12}=\frac{\theta r^2}{\theta r}=r\]

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

The thing is that you are using information that you already know to solve this but you just need to mess around with it.

OpenStudy (anonymous):

For example the area of a circle is pi * r^2 It can also rewrite it as 2* pi /2 * r^2 Where 2 * pi is 360 degrees or the full circle. But now we want the area of part of the circle so by knowing the degrees of the arc you can find a fraciton of it. For example if you wanted an arc that represents half of the circle it will be half of 360 or half of 2 * pi = pi so the area will be 1/2 * pi * r^2